OpenStudy (yumyum247):

Only math GODS are invited please!! Cheque my work plz!! Thanks you!! <3

OpenStudy (yanasidlinskiy):

I don't get what's the question? Unless I'm misreading it or something?

OpenStudy (yumyum247):

@jim_thompson5910

OpenStudy (yumyum247):

@ganeshie8

OpenStudy (yumyum247):

@phi

OpenStudy (phi):

1d. (a+b)^2 is (a+b)(a+b) that is not a^2 + b^2

OpenStudy (yumyum247):

wait plz let me check

OpenStudy (phi):

in the last line of 2a, you could factor out the -1 f(-x) = -(5x^7+6x) and then write f(-x) = -f(x) which is the definition of an odd function.

OpenStudy (yumyum247):

yes ser

OpenStudy (phi):

in 2c, it can be shown that sin(-x) is -sin(x) (if you look at a plot of sin x , it is odd) that means you can write f(-x) = -2 x sin(x) = f(x) which means you have an even function.

OpenStudy (yumyum247):

yes ser

OpenStudy (phi):

in 3) you can't cancel out logs you can use "rules" to combine the logs you should get 2 log_4 x

OpenStudy (yumyum247):

yes ser

OpenStudy (phi):

are you sure about h(x) = 2log_4 (2) that is a constant number

OpenStudy (yumyum247):

am not sure, tbh..

OpenStudy (phi):

where did it come from? you can simplify 2 log_4(2) to 2*1/2 = 1

OpenStudy (yumyum247):

shouldn't 2h(x) mean the function of h(x) should be multiplied by 2??

OpenStudy (phi):

yes. but you have \[ h(x) = 2\log_42\] there is no "x" on the right side. If you put the left side into wolfram http://www.wolframalpha.com/input/?i=2%5Clog_4(2) it will evaluate it: it's the same as h(x) =1 (though if you know 4^(1/2) = 2 then you know log_4(2) = 1/2 and 2*1/2 = 1)

OpenStudy (phi):

maybe it's a typo ? and you mean h(x) = 2 log_4(x) ?

OpenStudy (yumyum247):

no its 2log_4 2

OpenStudy (phi):

in that case, simplify it to h(x) = 1 (similar to y=1, a horizontal line if we plot it)

OpenStudy (phi):

your part 3) is a bit of a mess. but part 4) is good

OpenStudy (phi):

n 14c, you dropped a minus sign. the linear graph shows the starting number to be -797.82 the best start number is the number closest to 98 that is the exponential model (notice that that model has a curve that goes close to the plotted points.

OpenStudy (yumyum247):

you mean am suppose to drop the minus sign??

OpenStudy (yumyum247):

because -797.82 is not given in the tabel

OpenStudy (phi):

no, I mean you did drop the minus sign, and wrote The linear graph shows that the study of population of bacteria started with an initial number of 797.82 bacteria. and that is not correct

OpenStudy (yumyum247):

is -797.82 linear model doesn't give the best answer..????

OpenStudy (yumyum247):

then which model gives the best answer?!?!?

OpenStudy (phi):

The linear model predicts that there are -797.82 bacteria at time 0 not only is that far from truth (98), it does not make physicial sense) in other words, the linear model is a poor fit .

OpenStudy (yumyum247):

ok so than exponential function models the data best?!?!?

OpenStudy (yumyum247):

i mean the exponential gives the best answer?!?!?

OpenStudy (phi):

yes. it definitely matches the data in the table best. It has the most accurate prediction of the starting number (104 is close to 98, right ?)

OpenStudy (yumyum247):

ok the only thing that i need your help with is question numner 3c

OpenStudy (yumyum247):

i believe you sent you the actual question. can you please help me

OpenStudy (phi):

you should redo all of part 3) a , b, and c

OpenStudy (yumyum247):